# Collective Superradiance: Estimating the Peak Emission Rate and Time

### v1

@@ I Introduction
Quantum emitter ensembles interacting through a shared electromagnetic reservoir can exhibit cooperative behavior in the form of superradiance [1, 2, 3, 4, 5, 6, 7]. The case of initially fully inverted ensembles results in the cooperative enhancement of the radiative decay process and the arrival of a superradiant emission peak (see Fig. 1(b)). Accurately predicting the peak photon emission rate and time of such systems is not only a long-standing open problem in theoretical quantum optics but relevant in a wide range of research areas including superradiant lasing [8, 9, 10, 11], biochemistry [12], photochemistry [13, 14], radio astronomy [15, 16], nanomaterials [17], condensed matter physics [18], and quantum error correction [19, 20]. Superradiance in the few- and many-excitation regime has been observed numerous times over the years in a broad range of platforms such as atomic gases [21, 22, 23, 24, 25, 26], solid-state systems [27, 28, 29, 30, 31, 32], molecular emitters [33, 34], ensembles of nuclei [35] and waveguide platforms [27, 36], illustrated in Fig. 1(a). However, theoretically predicting the peak radiated power and time presents a significant computational challenge: tracking the time evolution of the density operator for an ensemble of $N$ quantum emitters, each undergoing collective spontaneous decay, quickly becomes intractable as $N$ increases due to the exponential growth in system dimension. The difficulty lies in the need to account for nonlinear all-to-all interactions between the two-level emitters, inherent in the superradiant decay process. Conventional methods that rely on direct simulation scale exponentially with $N$ [37, 38, 39], often requiring supercomputing resources for even moderately sized emitter numbers. In the simplest case of Dicke’s original superradiance formulation [1], all emitters are assumed to be indistinguishable, allowing even for analytic solutions [40]. For superradiance in extended ensembles [41, 42] numerical methods either rely on approximations in terms of emitter correlations [43, 44, 45], or on symmetries of the problem to reduce complexity [46], however generally the system’s dimension scales polynomially with the emitter number at best. Furthermore, variations in the system’s physical parameters, such as geometry, emitter spacing, and emitter types, add layers of complexity to the problem.
@@ Appendix B Quantum master equation and dipole-dipole interactions
where $\Gamma_{k}$ are the eigenvalues of the $N\times N$ dissipation matrix with elements $\Gamma_{mn}$ and $\mathcal{O}_{k}$ are the associated collective jump operators given by $\mathcal{O}_{k}=\sum_{n=1}^{N}c_{k,n}\sigma_{n}$ with coefficients $c_{k,n}$ fulfilling $\sum_{k=1}^{N}\Gamma_{k}|c_{k,n}|^{2}=\sum_{n}^{N}\Gamma_{n}/N$ and $\sum_{n=1}^{N}c^{*}_{k,n}c_{k^{\prime},n}=\delta_{kk^{\prime}}$. The coherent and dissipative dipole-dipole couplings between emitters $n$ and $m$ read

### v2

@@ I Introduction
Quantum emitter ensembles interacting through a shared electromagnetic reservoir can exhibit cooperative behavior in the form of superradiance [1, 2, 3, 4, 5, 6, 7]. The case of initially fully inverted ensembles results in the cooperative enhancement of the radiative decay process and the arrival of a superradiant emission peak (see Fig. 1(b)). Accurately predicting the peak photon emission rate and time of such systems is not only a long-standing open problem in theoretical quantum optics but relevant in a wide range of research areas including superradiant lasing [8, 9, 10, 11], biochemistry [12], photochemistry [13, 14], radio astronomy [15, 16], nanomaterials [17], condensed matter physics [18], and quantum error correction [19, 20]. Superradiance in the few- and many-excitation regime has been observed numerous times over the years in a broad range of platforms such as atomic gases [21, 22, 23, 24, 25, 26], solid-state systems [27, 28, 29, 30, 31, 32], molecular emitters [33, 34], ensembles of nuclei [35] and waveguide platforms [27, 36], illustrated in Fig. 1(a). However, theoretically predicting the peak radiated power and time presents a significant computational challenge: tracking the time evolution of the density operator for an ensemble of $N$ quantum emitters, each undergoing collective spontaneous decay, quickly becomes intractable as $N$ increases due to the exponential growth in system dimension. The difficulty lies in the need to account for nonlinear all-to-all interactions between the two-level emitters, inherent in the superradiant decay process. Conventional methods that rely on direct simulation scale exponentially with $N$ [37, 38, 39], often requiring supercomputing resources for even moderately sized emitter numbers. In the simplest case of Dicke’s original superradiance formulation [1], all emitters are assumed to be indistinguishable, allowing even for analytic solutions [40, 41]. For superradiance in extended ensembles [42, 43] numerical methods either rely on approximations in terms of emitter correlations [44, 45, 46], or on symmetries of the problem to reduce complexity [47], however generally the system’s dimension scales polynomially with the emitter number at best. Furthermore, variations in the system’s physical parameters, such as geometry, emitter spacing, and emitter types, add layers of complexity to the problem.
@@ Appendix B Quantum master equation and dipole-dipole interactions
where $\Gamma_{k}$ are the eigenvalues of the $N\times N$ dissipation matrix with elements $\Gamma_{mn}$ and $\mathcal{O}_{k}$ are the associated collective jump operators given by $\mathcal{O}_{k}=\sum_{n=1}^{N}c_{k,n}\sigma_{n}$ with coefficients $c_{k,n}$ fulfilling $\sum_{k=1}^{N}\Gamma_{k}|c_{k,n}|^{2}=\Gamma$ and $\sum_{n=1}^{N}c^{*}_{k,n}c_{k^{\prime},n}=\delta_{kk^{\prime}}$. The coherent and dissipative dipole-dipole couplings between emitters $n$ and $m$ read
@@ Appendix D Influence of the Hamiltonian on superradiant decay
Appendix E Partially excited ensemble in the Dicke limit
In this section, we briefly discuss the case of a partially excited ensemble of the form
$|\psi\rangle_{\varphi}=\bigotimes_{n=1}^{N}\Big{(}\sqrt{1-\varphi}|g_{n}\rangle+\sqrt{\varphi}|e_{n}\rangle\Big{)}=\sum_{m=1}^{N}\varphi^{m/2}(1-\varphi)^{(N-m)/2}\frac{(S^{\dagger})^{m}}{m!}\bigotimes_{n=1}^{N}|g_{n}\rangle,$ (19)
but for fully inverted ensembles with identical decay rates $\Gamma_{n}=\Gamma$, it is again given by Eq. 4 from the main text. The total excitation number at $t=0$ is now given by $N_{\alpha}\leq N$ and the peak emission by
$R_{\varphi}(0)=N\varphi^{2}+N^{2}\varphi(1-\varphi)$ (20)